# Refined climatologies of future precipitation over High Mountain Asia using probabilistic ensemble learning

### v1

@@ 3.2.2 Weights
We would like to choose weights $\bm{w}$ such that the MoE accurately predicts future and past precipitation distributions. Without prior knowledge, the simplest weighting would be $w_{r}=1/R$, i.e. an equally-weighted (EW) mixture of the RCM surrogates, not assigning more importance to any particular RCM.. Instead, we inform the weights through a reference dataset (APHRODITE; see Section 2.0.3), having previously downscaled RCM predictions onto the same grid using the surrogate models, as shown in Sections 3.1.1 and 4. A naive approach would be to directly learn the weights that maximise the likelihood of the data (APHRODITE). However, there are two potential pitfalls to this method. First, the number of effective parameters is $R-1$ and may result in overfitting. Second, direct maximum likelihood optimisation could upweigh RCM surrogates that, on their own, predict precipitation poorly, but whose sum best approximates APHRODITE. We want to ensure that RCM surrogates that contribute significantly to the aggregate are independently close fits to APHRODITE.
Thus, we inform the weights using a measure of distance between the distribution of each RCM surrogate and that of APHRODITE. We choose the Wasserstein distance to achieve this (Panaretos and Zemel, 2019). The Wasserstein distance originates from optimal transport theory and represents the minimum cost of moving the probability density of one distribution into the shape of another. For two empirical distributions $P$ and $Q$, the Wasserstein distance $W$ is given by:
$W(P,Q)=\left(\frac{1}{N}\sum^{N}_{i=1}\|X_{i}-Y_{i}\|^{p}\right)^{1/p}$ (10)
where $X_{i}$ and $Y_{i}$ are the samples of $P$ and $Q$, respectively, and $p$ is the moment of the distribution we are interested in. Here, we use the first moment $p=1$, which represents the shift of the distribution mean. We calculate Wasserstein distances with respect to a reference historical period for each APHRODITE grid point and month. Previous research has shown that evaluating the shapes of the distribution rather than the differences in the absolute precipitation outputs better distinguishes how well the models represent precipitation (Martinez-Villalobos and Neelin, 2021). We therefore scale the 95th percentile of the BCM outputs to match that of APHRODITE.
@@
$w_{r}=\frac{e^{-f(W_{r})/T}}{\sum_{r^{\prime}}e^{-f(W_{r^{\prime}})/T}},$ (11)
where $f(\cdot)$ is some monotonically increasing function. The statistical temperature $T$ is then optimised to maximise the likelihood on the APHRODITE data set. We chose $f(\cdot)=\ln(\cdot)$, which reduces the variation between the Wasserstein distances. This choice meets our criteria for the weights to smoothly vary in space and to not choose a single RCM. To further stabilise optimisation, we maximised the likelihood around a $5\times 5$ grid surrounding each location. The RCM surrogates with the smallest Wasserstein distances and the maximum MoE weights over HMA are plotted in C.
@@ 4 MoE validation
To evaluate the performance of the MoE approach relative to the EW, the historical experiment is divided into two sets: a training period (1951-1980) and a validation period (1981-2005). We use the continuous rank probability score (CRPS), which measures the distance between the predicted (MoE or EW) and observed (APHRODITE) cumulative distribution functions (CDF) $F$. The CRPS is defined as the integral of the squared difference between the predicted CDF $F(y)$ and the observed CDF $F_{\text{obs}}(y)$ over the range of possible values $y$:
@@
Figure 5: MoE and EW CRPS differences over the held-out validation period (1981-2005). Negative values (blue) imply MoE matches APHRODITE more closely while positive values (red) represent times and location where the EW performs better. The MoE outperforms the EW for each month, in particular over the Karakoram and Himalayan arc.
@@ 5 Refined climatologies
To generate refined historical and future climatologies, we optimise the weights over the entire historical period (1951-2005), thus maximising data availability. We then use these weights to generate MoE ensemble predictions for the same climate scenarios and time periods as Sanjay et al. (2017): i) historical from 1976 and 2005, ii) RCP4.5 from 2036 to 2065 (near future), and from 2066 to 2095 (far future), and iii) RCP8.5 from 2036 to 2065 (near future), and from 2066 to 2095 (far future). In the following analysis, we focus on historical and far-future periods to generate projected precipitation trends relative to a historical baseline. The near-future projections show trends comparable to the far future, but with less pronounced changes, and are included in E.
@@ 5.1 Historical predictions
Figure 6: Historical MoE and EW relative error for median precipitation (REMP) with respect to APHRODITE over HMA. The MoE REMP (top), the EW REMP (middle), and the REMP difference between MoE and EW (bottom) are plotted for the historical reference period (1976-2005) for the monsoon (left) and winter (right). The MoE makes large improvements over the EW, in particular, over locations where precipitation is overestimated by the EW.
Figure 6 plots the relative error for median precipitation (REMP) for the MoE and EW with respect to APHRODITE for the historical period. For the monsoon season, the MoE and EW both under- and overestimate precipitation for the Himalaya and Karakoram (HMA1 and HMA2). More specifically, the models predict a wet bias for the Karakoram, inner Himalayas, and western Tibetan Plateau and a dry bias for the outer Himalayas. Precipitation over the eastern Tibetan Plateau, Hengduan Shan, and Southeast Tibet (HMA3) is generally underestimated by both models. For winter, the MoE under- and overestimate precipitation across the three studied subregions with greater wet biases, including the Tibetan Plateau and Hengduan Shan. On the other hand, the EW scheme largely overestimates precipitation across all three areas during winter.
@@ 5.2 Future projections
Figure 7: Relative changes between far-future (2066-2095) and historical (1976-2005) MoE precipitation. The plot shows the median changes for RCP4.5 (top) and RCP8.5 (bottom) over the monsoon (left) and winter (right) for the far-future across HMA. The MoE predicts large departures from the precipitation predicted for the historical reference period.
@@
Figure 8: MoE and EW prediction differences for the far future (2066-2095) under RCP8.5 across HMA. The plot shows the differences between the predicted relative changes for the 5th percentile (bottom), median (middle), and 95th (top) during the monsoon (left) and winter (right). Although for many locations the MoE and EW projections are similar, there are significant difference in predictions during both the monsoon and winter over key mountain ranges.
@@ 6 Discussion
This paper has explored a probabilistic ensemble learning approach to aggregating RCMs using a MoE approach, with the aim of improving future precipitation projections over using the EW. Unlike other ensembling techniques, our method does not narrow the non-discountable envelope of climate change, but refines it. Although not shown, the method should work equally well with RCMs on different grids and can be straightforwardly applied to any kind of model ensemble. However, this study does present several limitations.
First, we assume that precipitation from the driving GCMs is sufficiently well-modelled to dynamically downscaled. In reality, GCMs present large differences in their precipitation predictions over HMA (Panday et al., 2015, Palazzi et al., 2015). Moreover, recent studies incorporating more realistic aerosol scenarios further amplify future precipitation uncertainty (Jiang et al., 2023, Wilcox et al., 2020). The poor representation of precipitation in GCMs can be observed in this study through the optimal weights chosen by the MoE: when no model in the ensemble performs particularly well, the MoE chooses weights close to the EW. This is the case for the Hindu Kush during winter and for the Tibetan Plateau during summer (cf. Figure 6 and Figure 10). This behaviour, in turn, implies poor representation of two main precipitation drivers for HMA: the East Asian monsoon and western disturbances. Further work could use this framework to explain performance differences across RCMs by analysing other output variables, such as pressure or humidity fields.
@@ 7 Conclusion
This paper investigated a mixture of experts (MoE) approach to aggregate an ensemble of RCMs. This ensemble learning method is generalisable to other climate variables and model ensembles and could be applied to model aggregation problems in other fields. The MoE was applied to the CORDEX-WAS ensemble over HMA and was found to be more accurate than the equally-weighted average (EW) or any single ensemble member for a held-out validation period. Compared to the EW, the MoE projects a significantly wetter summer but a drier winter over the western Himalayas and Karakoram and a significantly wetter winter over the Tibetan Plateau, South East Tibet and Hengduan Shan for both the near and far future under RCP8.5 and RCP4.5.
@@
The code for this paper is available on GitHub (https://github.com/kenzaxtazi/bcm4rcm). The CORDEX-WAS data are available through the Earth System Federation Grid nodes https://esgf-metagrid.cloud.dkrz.de and the APHRODITE data through the APHRODITE’s Water Resources website http://aphrodite.st.hirosaki-u.ac.jp/download/.
@@ Acknowledgements
This work was supported by the Engineering and Physical Sciences Research Council through the AI for Environmental Risk (AI4ER) CDT [grant number: EP/S022961/1] for KT and MGM, and the DTP International Studentship [grant number EP/W524475/1] for SWPK.
@@ Appendix C Which RCMs, when and where?
Figure 10: Maximum MoE weights for each month and location over HMA. The weights are optimised between 1951 an 2005. Here, one RCM surrogate (BCM) is never chosen over all others; the maximum $w_{r}$ is never 1. The maximum BCM weighting is highest over the Hindu Kush during the summer and over the Tibetan Plateau during the winter. Conversely, the MoE with close to equal-weights (i.e. maximum $w_{r}\rightarrow 1/13$) is more advantageous during the monsoon over the Tibetan Plateau and during the winter over Hindu Kush.
Appendix D MoE and RCM annual averaged CRPS differences
Figure 11: Annual CRPS difference between the MoE and RCM surrogates (BCM) for the held-out validation period (1981-2005). Negative values (blue) imply MoE matches APHRODITE more closely while positive values (red) represent times and location where the BCM performs better. Overall, the MoE outperforms the BCMs over the entire year with large improvement over HM1 and HM2, especially for the RegCM4 surrogates.
Figure 12: Scorecard showing the differences between the CRPS yielded by the MoE and the CRPS yielded by each RCM surrogate (BCM) as well as the EW over the held out validation period (1981-2005). Each cell represents the difference for a specific model and month, averaged over the entire spatial domain. Positive values (red) indicate higher CRPS for the MoE compared to the BCM, while negative values (blue) indicate lower CRPS for the MoE compared to the BCM.
@@ Appendix E Near-future predictions
Figure 13: Relative changes between near-future (2036-2065) and historical (1976-2005) MoE precipitation across HMA. The plot shows the median changes for RCP4.5 (top) and RCP8.5 (bottom) during the monsoon (left) and winter (right). Results are similar to far-future predictions with less pronounced changes.
Figure 14: MoE and EW prediction differences for the near future (2066-2095) under RCP8.5 across HMA. The plot shows the difference between the predicted relative changes for the 5th percentile (bottom), median (middle), and 95th(top) during the monsoon (left) and winter (right). Results are similar to far-future predictions with less pronounced differences.

### v2

@@ 3.2.2 Weights
We would like to choose weights $\bm{w}$ such that the MoE accurately predicts future and past precipitation distributions. Without prior knowledge, the simplest weighting would be $w_{r}=1/R$, i.e. an equally-weighted (EW) mixture of the RCM surrogates, not assigning more importance to any particular RCM. Instead, we inform the weights through a reference dataset (APHRODITE; see Section 2.0.3), having previously downscaled RCM predictions onto the same grid using the surrogate models, as shown in Sections 3.1.1 and 4. A naive approach would be to directly learn the weights that maximise the likelihood of the parameters (APHRODITE). However, there are two potential pitfalls to this method. First, the number of effective parameters is $R-1$ (=12), compared to the number of datapoints at each location and time (${\sim}30$), resulted in overfitting the $R$ model weights (not shown). Second, direct maximum likelihood optimisation could upweigh RCM surrogates that, individually predict precipitation poorly, but when summed together best approximate APHRODITE. We want to ensure that RCM surrogates that contribute significantly to the aggregate are independently close fits to APHRODITE.
Thus, we inform the weights using a measure of distance between the distribution of each RCM surrogate and that of APHRODITE. We choose the Wasserstein distance (Panaretos and Zemel, 2019), which originates from optimal transport theory and represents the minimum cost of moving the probability density of one distribution into the shape of another. For two empirical distributions $P$ and $Q$, the Wasserstein distance $W$ is given by:
$W(P,Q)=\left(\frac{1}{N}\sum^{N}_{i=1}\|X_{i}-Y_{i}\|^{p}\right)^{1/p},$ (10)
where $X_{i}$ and $Y_{i}$ are the samples of $P$ and $Q$, respectively, and $p$ is the moment of the distribution we are interested in. Here, we use the first moment $p=1$, which represents the shift of the distribution mean. We calculate the Wasserstein distances with respect to a reference historical period for each APHRODITE grid point and month. Previous research has shown that evaluating the shapes of the distribution rather than the differences in the absolute precipitation outputs better distinguishes how well the models represent precipitation (Martinez-Villalobos and Neelin, 2021). We therefore scale the 95th percentile of the BCM outputs to match that of APHRODITE.
@@
$w_{r}=\frac{e^{-h(W_{r})/T}}{\sum_{r^{\prime}}e^{-h(W_{r^{\prime}})/T}},$ (11)
where $h(\cdot)$ is some monotonically increasing function. The statistical temperature $T$ is then optimised to maximise the likelihood on the APHRODITE dataset. We chose $h(\cdot)=\ln(\cdot)$, which reduces the variation between the Wasserstein distances. This choice meets our criteria for the weights to vary smoothly in space and to not choose a single RCM. To further stabilise optimisation, we maximised the likelihood around a $5\times 5$ grid surrounding each location. The RCM surrogates with the smallest Wasserstein distances and the maximum MoE weights over HMA are shown in C.
@@ 4 MoE validation
To evaluate the performance of the MoE approach relative to the EW, the historical experiment is divided into two sets: a training period (1951–1980) and a validation period (1981–2005). We use the continuous rank probability score (CRPS, zamo2018estimation), which measures the distance between the predicted (MoE or EW) and observed (APHRODITE) cumulative distribution functions (CDF) $F$. The CRPS is defined as the integral of the squared difference between the predicted CDF $F(y)$ and the observed CDF $F_{\text{obs}}(y)$ over the range of possible values $y$:
@@
Figure 5: MoE and EW CRPS differences over the held-out validation period (1981–2005). Negative values (blue) imply MoE matches APHRODITE more closely while positive values (red) represent times and location where the EW performs better. The MoE outperforms the EW for each month, in particular over the Karakoram and Himalayan arc.
@@ 5 Refined climatologies
To generate refined historical and future climatologies, we optimise the weights over the entire historical period (1951–2005), thus maximising data availability. We then use these weights to generate MoE ensemble predictions for the same climate scenarios and time periods as Sanjay et al. (2017): i) historical (1976–2005), ii) RCP4.5 near (2036–2065) and future (2066–2095), and iii) RCP8.5 near and far future. In the following analysis, we focus on historical and far-future periods to generate projected precipitation trends relative to a historical baseline. The near-future projections show trends comparable to the far future, but with less pronounced changes, and are included in E.
@@ 5.1 Historical predictions
Figure 6: Historical MoE and EW relative error for median precipitation (REMP) with respect to APHRODITE over HMA. The MoE REMP (top), the EW REMP (middle), and the REMP difference between MoE and EW (bottom) are plotted for the historical reference period (1976–2005) for the monsoon (left) and winter (right). The MoE makes large improvements over the EW, in particular, over locations where precipitation is overestimated by the EW.
Figure 6 plots the relative error for median precipitation (REMP) for the MoE and EW with respect to APHRODITE for the historical period. For the monsoon season, the MoE and EW both under- and overestimate precipitation for the Himalaya and Karakoram (HMA1 and HMA2). More specifically, the models predict a wet bias for the Karakoram, inner Himalayas, and western Tibetan Plateau and a dry bias for the outer Himalayas. Precipitation over the eastern Tibetan Plateau, Hengduan Shan, and Southeast Tibet (HMA3) is generally underestimated by both models. For winter, the MoE under- and overestimates precipitation across the three studied subregions with greater wet biases, including the Tibetan Plateau and Hengduan Shan. On the other hand, the EW scheme largely overestimates precipitation across all three areas during winter.
@@ 5.2 Future projections
Figure 7: Relative changes between far-future (2066–2095) and historical (1976–2005) MoE precipitation. The plot shows the median changes for RCP4.5 (top) and RCP8.5 (bottom) over the monsoon (left) and winter (right) for the far-future across HMA. The MoE predicts large departures from the precipitation predicted for the historical reference period.
@@
Figure 8: MoE and EW prediction differences for the far future (2066–2095) under RCP8.5 across HMA. The plot shows the differences between the predicted relative changes for the 5th percentile (bottom), median (middle), and 95th (top) during the monsoon (left) and winter (right). Although for many locations the MoE and EW projections are similar, there are significant difference in predictions during both the monsoon and winter over key mountain ranges.
@@ 6 Discussion
This paper has explored a probabilistic ensemble learning approach to aggregating RCMs using a MoE approach, with the aim of improving future precipitation projections over using the EW. Unlike other ensembling techniques, our method does not narrow the non-discountable envelope of climate change, but refines it. With GPs as its foundational building blocks, this method should work equally well with RCMs on different grids and can be straightforwardly applied to any kind of model ensemble. However, this study does present several limitations.
First, we assume that precipitation from the driving GCMs is sufficiently well-modelled to be dynamically downscaled. In reality, CMIP5 GCMs present large differences in their precipitation predictions over HMA (Panday et al., 2015, Palazzi et al., 2015). Moreover, recent studies incorporating more realistic aerosol scenarios further amplify future precipitation uncertainty (Jiang et al., 2023, Wilcox et al., 2020). The poor representation of precipitation in GCMs can be observed in this study through the optimal weights chosen by the MoE: when no model in the ensemble performs particularly well, the MoE chooses weights close to the EW. This is the case for the Hindu Kush during winter and for the Tibetan Plateau during summer (cf. Figure 6 and Figure 10). This behaviour, in turn, implies poor representation of two main precipitation drivers for HMA: the East Asian summer monsoon and western disturbances. Further work could use this framework to explain performance differences across RCMs by analysing other output variables, such as pressure or humidity fields.
@@ 7 Conclusion
This paper investigated a mixture of experts (MoE) approach to aggregate an ensemble of RCMs. This ensemble learning method is generalisable to other climate variables and model ensembles and could be applied to model aggregation problems in other fields. The MoE was applied to the CORDEX-WAS ensemble over HMA and was found to be more accurate than the equally-weighted average (EW) or any single ensemble member for a held-out validation period. Compared to the EW, the MoE projects a significantly wetter summer but a drier winter over the western Himalayas and Karakoram and a significantly wetter winter over the Tibetan Plateau, Hengduan Shan, and South East Tibet for both the near and far future under RCP4.5 and RCP8.5.
@@
The code and data for this paper are available on GitHub (https://github.com/kenzaxtazi/bcm4rcm). The CORDEX-WAS data are available through the Earth System Federation Grid nodes https://esgf-metagrid.cloud.dkrz.de and the APHRODITE data through the APHRODITE’s Water Resources website http://aphrodite.st.hirosaki-u.ac.jp/download/.
@@ Acknowledgements
The authors thank Andrew Orr and Scott J. Hosking for their thoughtful feedback on the manuscript. This work was supported by the Engineering and Physical Sciences Research Council through the AI for Environmental Risk (AI4ER) CDT [grant number: EP/S022961/1] for KT and MGM, and the DTP International Studentship [grant number EP/W524475/1] for SWPK.
@@ Appendix C Which RCMs, when and where?
Figure 10: Maximum MoE weights for each month and location over HMA. The weights are optimised between 1951 and 2005. Here, one RCM surrogate (BCM) is never chosen over all others; the maximum $w_{r}$ is never 1. The maximum BCM weighting is highest over the Hindu Kush during the summer and over the Tibetan Plateau during the winter. Conversely, the MoE with close to equal-weights (i.e. maximum $w_{r}\rightarrow 1/13$) is more advantageous during the monsoon over the Tibetan Plateau and during the winter over Hindu Kush.
Appendix D MoE and BCM CRPS differences
Figure 11: Annual CRPS difference between the MoE and RCM surrogates (BCM) for the held-out validation period (1981–2005). Negative values (blue) imply MoE matches APHRODITE more closely while positive values (red) represent times and location where the BCM performs better. Overall, the MoE outperforms the BCMs over the entire year with large improvement over HM1 and HM2, especially for the RegCM4 surrogates.
Figure 12: Scorecard showing the differences between the CRPS yielded by the MoE and the CRPS yielded by each RCM surrogate (BCM) as well as the EW over the held out validation period (1981–2005). Each cell represents the difference for a specific model and month, averaged over the entire spatial domain. Positive values (red) indicate higher CRPS for the MoE compared to the BCM, while negative values (blue) indicate lower CRPS for the MoE compared to the BCM.
@@ Appendix E Near-future predictions
Figure 13: Relative changes between near-future (2036–2065) and historical (1976–2005) MoE precipitation across HMA. The plot shows the median changes for RCP4.5 (top) and RCP8.5 (bottom) during the monsoon (left) and winter (right). Results are similar to far-future predictions with less pronounced changes.
Figure 14: MoE and EW prediction differences for the near future (2066–2095) under RCP8.5 across HMA. The plot shows the difference between the predicted relative changes for the 5th percentile (bottom), median (middle), and 95th(top) during the monsoon (left) and winter (right). Results are similar to far-future predictions with less pronounced differences.